I have a geometry question I need help with it is as follows:
A triangle has the following sides: 12 feet, 20 feet, and 16 feet. Is this a right triangle? Explain your answer.

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

A triangle has the following sides: 12 feet, 20 feet, and 16 feet.
-----
Yes, if 20^2 = 12^2 + 16^2
It does.
Cheers,
Stan H.

Answer by richwmiller(17219)

let's put them in order of size
12,16,20
divide by 4
3,4,5 yes it is the classic 3,4,5 right triangle
What is another way using the Pythagorean theorem ?
the largest side will be the hypotenuse
12^2+16^2=20^2
144+16^2=400
Does 16^2=256?
if so then it is a right triangle.